# Monopole Acceleration in Intergalactic Magnetic Fields

### v1

@@ 4 Conclusion
We also showed that if the cosmic flux of MMs is larger than the threshold values in Eqs. (2.17) or (2.18) (which are also illustrated in Figure 1), their backreaction on IGMFs is non-negligible. In this case the energy oscillates between the IGMFs and the MMs, and the oscillation-averaged MM velocity depends also on the MM density. Future developments in studies of IGMFs may be able to constrain such IGMF-MM oscillations, which will allow us to improve bounds on the MM flux.
This paper serves as a first step towards a complete understanding of MM acceleration in cosmic environments. An accurate knowledge of the velocity of MMs entering the Earth is crucial for interpreting the constraints from terrestrial detectors (e.g. [27, 28]) in terms of the MM mass. We will discuss the implications of acceleration effects for MM experiments in a companion work [29]. It would also be interesting to study relativistic MMs as a possible explanation of the recently claimed detection of an extremely high-energy cosmic ray [30]; see also the earlier papers [31, 14, 15, 32, 33] which studied similar ideas. Finally, we note that a cosmic population of relativistic MMs from IGMFs can serve as dark radiation, which may be further constrained from cosmological studies.
@@ Acknowledgments
We thank Ivan De Mitri, Pranjal Ralegankar, Piero Ullio, and Matteo Viel for helpful discussions. T.K. acknowledges support from INFN TAsP, PRIN MIUR (code 20224JR28W), and JSPS KAKENHI (Grant No. JP22K03595).
Appendix A Energy loss of fast monopoles in a nonrelativistic plasma
@@
$\frac{d\bm{p}}{dt}=\frac{qg}{4\pi}\frac{\bm{u}\times\bm{r}}{r^{3}}.$ (A.1)
@@
$\frac{d\sigma}{d\theta}=-2\pi b\frac{db}{d\theta}=\frac{(qg)^{2}}{2\pi p^{2}\theta^{3}}.$ (A.3)
We assume that in the rest frame of the plasma, the MM velocity $\bm{v}$ has a much larger amplitude compared to the velocities of the individual plasma particles. [^8] A large relative velocity between the MM and the plasma particles also supports the procedure of focusing on small-angle scatterings. Then in the MM rest frame, the particles can be considered to move with a universal velocity $\bm{u}=-\bm{v}$, and the scatterings with a collection of the particles induce the MM to experience a force,
@@
The expression for the drag force can be covariantized by noting that the force vector $\mathcal{F}^{\mu}$ should, in the MM rest frame, have spatial components matching with (A.4), along with a vanishing time component since the MM’s energy does not change. These are satisfied by a force vector with the form,
$\mathcal{F}^{\mu}=-\frac{C(qg)^{2}n_{\mathrm{p}}}{4\pi m_{\mathrm{p}}}\frac{1}{v_{\mathrm{rel}}}\left\{\frac{\mathrm{u}^{\mu}}{\mathrm{u}_{\nu}\mathrm{v}^{\nu}}+\mathrm{v}^{\mu}\right\},$ (A.6)
@@
$m\mathrm{v}^{\nu}\nabla_{\nu}\mathrm{v}^{\mu}=g\tilde{F}^{\mu\nu}\mathrm{v}_{\nu}+\mathcal{F}^{\mu},$ (A.7)
@@
$m\frac{d}{dt}(\gamma\bm{v})=g\bm{B}-\frac{C(qg)^{2}n_{\mathrm{p}}}{4\pi m_{\mathrm{p}}}\frac{\bm{v}}{v}.$ (A.8)
Let us now compare the deceleration and acceleration terms in (A.8) in cosmic environments, by estimating their ratio:
$\epsilon=\frac{1}{gB}\frac{C(qg)^{2}n_{\mathrm{p}}}{4\pi m_{\mathrm{p}}}.$ (A.9)
@@
Let us comment on effects that we have not taken into account. Firstly, the above calculation is based on classical single-particle scatterings. A quantum field theory computation may give rise to corrections to the scattering cross section, especially in the ultrarelativistic regime. We also note that the picture of single-particle scatterings breaks down at distances larger than the Debye length, where plasma effects become important. These have been studied in [35, 36] for slow MMs, for which it was shown that the energy loss from plasma effects is typically not significantly larger than that from single-particle scatterings. Finally, we remark that MMs can further loose energy through ionization of neutral particles, atomic excitations, bremsstrahlung, electron-pair production, and photonuclear interactions; the last two processes become particularly important in the ultrarelativistic regime. (See e.g. [16] for a comprehensive study of energy-loss processes for MMs. A rough estimate of the interactions with neutral particles in the ISM is given in [25], according to which they are negligible for a wide range of parameters.) It would be important to study all these effects, in particular in the ultrarelativistic regime. We leave this for future work.
@@
$P\sim g^{2}\gamma^{6}\left\{(\dot{\bm{v}})^{2}-(\bm{v}\times\dot{\bm{v}})^{2}\right\}=\frac{g^{4}B^{2}}{m^{2}},$ (A.10)
where an overdot denotes a time-derivative. Upon moving to the far right-hand side, we have considered a MM accelerating along a homogeneous magnetic field and dropped the cross product, and also used the equation of motion (A.8) without the drag force term. The ratio between the radiative energy loss, and energy gain per time from the acceleration in the magnetic field, $gBv$ (here we take $g>0$), is thus
$\delta\sim\frac{g^{3}B}{m^{2}v}.$ (A.11)
@@
Appendix B Monopole-magnetic field oscillation
If MMs accelerated in magnetic fields do not dissipate their kinetic energy into the ambient plasma, they eventually return the energy to the magnetic fields. In this way the energy can move back and forth between the magnetic field and monopoles. However one would expect such oscillations to be subject to Landau damping, if the phase velocity of the magnetic field were smaller than the random velocity of the individual MMs [18].
In order to estimate the phase velocity, we start by considering a homogeneous one-dimensional system where the magnetic field has a component $B$ along a certain direction, and the MMs with charge $g$ ($>0$) have a velocity component $v$ along the same direction. (In this appendix, $B$ and $v$ are the not the amplitudes and thus can be negative.) Then by combining the MMs’ equation of motion, $m\,d(\gamma v)/dt=gB$, with the conservation of total energy density $\rho_{\mathrm{tot}}=nm\gamma+B^{2}/2$ where $n$ is the MM number density, one obtains an evolution equation for the magnetic field,
$\frac{d^{2}B}{dt^{2}}+\Omega^{2}B=0.$ (B.1)
@@ Appendix B Monopole-magnetic field oscillation
$\Omega=\frac{g}{\gamma^{3/2}}\left(\frac{n}{m}\right)^{1/2}.$ (B.2)
@@
$v_{\mathrm{ph}}\sim 10^{-5}\left(\frac{v_{\mathrm{MW}}}{10^{-3}}\right)^{-1/2}\left(\frac{F_{\mathrm{MW}}}{10^{-15}\,\mathrm{cm}^{-2}\mathrm{sec}^{-1}\mathrm{sr}^{-1}}\right)^{1/2}\left(\frac{g}{g_{\mathrm{D}}}\right)\left(\frac{m}{10^{18}\,\mathrm{GeV}}\right)^{-1/2}\left(\frac{\lambda_{\mathrm{G}}}{1\,\mathrm{kpc}}\right).$ (B.3)
We have rewritten $n$ in terms of the MM flux in the rest frame of the Milky Way Galaxy (cf. below (3.2); however here we consider both clustered and unclustered monopoles), and set the reference value for the flux to the Galactic Parker limit at $m=10^{18}\,\mathrm{GeV}$ (cf. Figure 4). Note that MMs with a Dirac charge and mass $m\gtrsim 10^{18}\,\mathrm{GeV}$ can cluster with the Galaxy [18, 25]. Such clustered MMs obtain a virial velocity of $10^{-3}$, which is larger than the magnetic field’s phase velocity; hence the oscillation is expected to rapidly evaporate. The phase velocity can in principle become larger for smaller masses, however light MMs do not cluster and hence should pass through the Galaxy before (completely) returning the energy to the Galactic fields. A flux much larger than the Parker limit can also increase the phase velocity beyond the virial velocity, however [37] claimed that oscillating Galactic fields do not match with observations, and also that the inhomogeneity in the MM distribution further leads to the damping of the oscillations. For these reasons, one can derive the Parker bound without worrying about the possibility that the Galactic fields survive as oscillating fields. [^10] Primordial magnetic fields can also be used to obtain Parker-type bounds, because MMs effectively dissipate energy into the primordial plasma and thus energy oscillations do not happen even without Landau damping [23, 24, 25]. This is in contrast to the minimal energy loss of MMs in the present-day IGM or ISM, which we showed in Appendix A.
@@
$v_{\mathrm{ph}}\sim 10^{-5}\,v_{\mathrm{CMB}}^{-1/2}\left(\frac{F_{\mathrm{CMB}}}{10^{-27}\,\mathrm{cm}^{-2}\mathrm{sec}^{-1}\mathrm{sr}^{-1}}\right)^{1/2}\left(\frac{g}{g_{\mathrm{D}}}\right)\left(\frac{m}{10^{10}\,\mathrm{GeV}}\right)^{-1/2}\left(\frac{\lambda_{\mathrm{I}}}{1\,\mathrm{Mpc}}\right).$ (B.4)
@@
Appendix C Monopole acceleration in intergalactic and galactic magnetic fields
MMs that cross the Earth are accelerated both by intergalactic and galactic magnetic fields. Depending on the amplitude and coherence length of IGMFs, and the MM number density and mass, one of the contributions is larger than the other. In this appendix we show the expected MM velocity at the Earth for different values of the MM mass and flux at the Earth, assuming a simple model for MM acceleration in Galactic fields.
Assuming Galactic magnetic fields with coherence length $\lambda_{\mathrm{G}}$, the velocities acquired by the MMs accelerated in Galactic magnetic fields is approximately:
$\left(\gamma v\right)_{\mathrm{G}}\sim\frac{gB_{\mathrm{G}}\lambda_{\mathrm{G}}}{m}.$ (C.1)
Therefore, the velocity of MMs on the Earth can be expressed as:
$v_{\mathrm{Earth}}=\mathrm{max.}\left\{v_{\mathrm{G}},v_{\mathrm{MW}}\right\},$ (C.2)
where $v_{\mathrm{MW}}$ is given by the expression in Eq. (3.2).
In Figure 5 we show the value of $v_{\mathrm{Earth}}$ in Eq. (C.2) as a function of the MM mass and flux at the Earth $F_{\mathrm{Earth}}=F_{\mathrm{MW}}$. In the figure we assume homogeneous IGMFs with $B=10^{-10}~{}\mathrm{G}$, and Galactic magnetic fields with amplitude $B_{\mathrm{G}}=10^{-6}~{}\mathrm{G}$ and $\lambda_{\mathrm{G}}=1~{}\mathrm{kpc}$. We also assume for the MM charge $g=g_{\mathrm{D}}$. We use the same color scheme of Figure 2. The plot shows three different behaviors of the MM velocity. In the upper region, the acceleration from Galactic fields is dominant over the acceleration from IGMFs. In this case, the value of the velocity is set by $v_{\mathrm{G}}$ and does not depend on the MM flux. In the central region, the velocity is set by $v_{\mathrm{MW}}$ and the backreaction on the IGMFs is non-negligible. In this case, the velocity depends on the MM flux in the CMB rest frame, and $F_{\mathrm{CMB}}=F_{\mathrm{Earth}}$. In the lower region, the velocity is still set by $v_{\mathrm{MW}}$ while the backreaction is negligible. Therefore, the velocity is independent of the MM flux. On the right part of the plot, the velocity is set by the peculiar velocity of the Milky Way, $v_{\mathrm{p}}\sim 10^{-3}$. The region excluded by the Galactic Parker bound is shown in gray.

### v2

@@ 4 Conclusion
Any additional magnetic fields around the Milky Way, such as those transported by galactic winds or outflow-driven bubbles, can also accelerate the MMs and further weaken the Parker bound. An accurate study of the effects of such fields requires numerical analyses, which we leave for future work.
We also showed that if the cosmic flux of MMs is larger than the threshold values of Eqs. (2.17) or (2.18) (which are also illustrated in Figure 1), their backreaction on IGMFs is non-negligible. In this case the energy oscillates between the IGMFs and the MMs, and the oscillation-averaged MM velocity depends on the MM density. Future developments in studies of IGMFs may be able to constrain such IGMF-MM oscillations, which will allow us to improve bounds on the MM flux.
This paper serves as a first step towards a complete understanding of MM acceleration in cosmic environments. An accurate knowledge of the velocity of MMs entering the Earth is crucial for interpreting the constraints from terrestrial detectors (e.g. [37, 38]) in terms of the MM mass. The acceleration in IGMFs can outweigh the acceleration by Galactic fields, as we showed in Appendix D. We will discuss the implications for MM experiments in a companion work [39]. It would also be interesting to study relativistic MMs as a possible explanation of the recently claimed detection of an extremely high-energy cosmic ray [40]; see also the earlier papers [41, 16, 17, 42, 43] which studied similar ideas. Finally, we note that a cosmic population of relativistic MMs from IGMFs can serve as dark radiation, which may be further constrained from cosmological studies.
@@ Acknowledgments
We thank Ivan De Mitri, Pranjal Ralegankar, Piero Ullio, and Matteo Viel for helpful discussions. The work of T.K. was supported by the European Union - NextGenerationEU, in the framework of the PRIN Project “Charting unexplored avenues in Dark Matter” (20224JR28W). T.K. also acknowledges support from INFN TAsP and JSPS KAKENHI (JP22K03595). MD acknowledges support by the Italian MUR Departments of Excellence grant 2023-2027 “Quantum Frontiers.”
Appendix A Primordial contribution to the monopole velocity
In the main text, we did not specify the origin of the IGMFs. In particular in Section 2, we evaluated the velocity of MMs that have been accelerated in IGMFs for a Hubble time. However it is often claimed in the literature that IGMFs have a primordial origin. [^8] See, e.g., [44] for an alternative explanation. Based on this premise, one might inquire about the relevance of MM acceleration in primordial magnetic fields, since primordial fields would have been much stronger at higher redshifts. In this appendix, we demonstrate that the acceleration in primordial fields only give an order-unity correction to the results derived in Section 2.
Let us consider for simplicity a homogeneous primordial magnetic field. The equation of motion of a MM in an expanding universe filled with a homogeneous magnetic field with strength $B$ is [33, 34]
$m\frac{d}{dt}(\gamma v)=gB-f_{\mathrm{p}}v-mH\gamma v,$ (A.1)
where we ignored velocity components perpendicular to the direction of the magnetic field. The last two terms in the right-hand side denote frictional forces from the MM scattering on charged particles in the primordial plasma, and from the cosmological expansion, with $H$ being the Hubble rate.
Before electron-positron annihilation, one can write the friction coefficient $f_{\mathrm{p}}$ as [32, 33]
$f_{p}\sim\frac{e^{2}g^{2}\mathcal{N}_{c}}{16\pi^{2}}T^{2},$ (A.2)
where $\mathcal{N}_{c}$ is the number of relativistic and electrically charged degrees of freedom in thermal equilibrium, and $T$ is the temperature of the plasma. Considering MMs with charges not too far from the Dirac charge, the Hubble friction is negligible compared to the friction from the plasma [33, 34]. The MMs thus approach a terminal velocity which results from equating the first two terms in the right-hand side of Eq. (A.1) as
$v_{\text{p}}=\frac{gB}{f_{\mathrm{p}}}\sim 10^{-8}\left(\frac{g_{D}}{g}\right)\left(\frac{B_{\rm I}}{10^{-15}G}\right).$ (A.3)
We note that the terminal velocity before the electron-positron annihilation is constant, since the homogeneous magnetic field $B$ and the friction coefficient $f_{\mathrm{p}}\propto T^{2}$ both redshift [^9] Magnetic fields with small coherence lengths can evolve differently, when the wave mode is inside the Hubble horizon [10]. as $(1+z)^{2}$. In the far right-hand side of Eq. (A.3), $B_{\mathrm{I}}$ represents the present-day magnetic field strength.
After electron-positron annihilation, the number of the free charged particles drops by $10$ orders of magnitude, so one can neglect the plasma friction, i.e. the second term in the right-hand side of Eq. (A.1). Then the present-day MM velocity can be obtained by integrating Eq. (A.1) as
[…]
where $z_{\text{an}}\sim 10^{9}$ is the redshift at electron-positron annihilation. Since the homogeneous magnetic field redshifts as $B(z)\propto(1+z)^{2}$, the only redshift dependence in the integrand comes from the Hubble rate $H(z)$. Upon moving to the far right-hand side, we have used the cosmological parameters for $\Lambda$CDM cosmology [45] and numerically integrated the second term. Using Eq. (A.3), the ratio between the two terms in Eq. (A.4) is written as
$\frac{(\gamma v)_{\rm p}}{1+z_{\rm an}}\left(\frac{gB_{\mathrm{I}}}{mH_{0}}\right)^{-1}\sim 10^{-8}\left(\frac{g_{\mathrm{D}}}{g}\right)^{2}\left(\frac{m}{10^{18}\,\mathrm{GeV}}\right),$ (A.5)
which shows that the first term is negligible for MM masses considered in this paper. Hence we obtain the velocity of MMs accelerated in a homogeneous primordial magnetic field as
$(\gamma v)_{0}\approx 3\times\frac{gB_{\mathrm{I}}}{mH_{0}}.$ (A.6)
This matches with the result given in Eq. (2.10) at the order-of-magnitude level. We therefore conclude that the order-of-magnitude results of this paper remain unaffected, even if the IGMFs have a primordial origin.
Appendix B Energy loss of fast monopoles in a nonrelativistic plasma
@@
$\frac{d\bm{p}}{dt}=\frac{qg}{4\pi}\frac{\bm{u}\times\bm{r}}{r^{3}}.$ (B.1)
@@
$\frac{d\sigma}{d\theta}=-2\pi b\frac{db}{d\theta}=\frac{(qg)^{2}}{2\pi p^{2}\theta^{3}}.$ (B.3)
We assume that in the rest frame of the plasma, the MM velocity $\bm{v}$ has a much larger amplitude compared to the velocities of the individual plasma particles. [^10] A large relative velocity between the MM and the plasma particles also supports the assumption of an unperturbed particle trajectory. Then in the MM rest frame, the particles can be considered to move with a universal velocity $\bm{u}=-\bm{v}$, and the scatterings with a collection of the particles induce the MM to experience a force,
@@
The expression for the drag force can be covariantized by noting that the force vector $\mathcal{F}^{\mu}$ should, in the MM rest frame, have spatial components matching with (B.4), along with a vanishing time component since the MM’s energy does not change. These are satisfied by a force vector with the form,
$\mathcal{F}^{\mu}=-\frac{C(qg)^{2}n_{\mathrm{p}}}{4\pi m_{\mathrm{p}}}\frac{1}{v_{\mathrm{rel}}}\left\{\frac{\mathrm{u}^{\mu}}{\mathrm{u}_{\nu}\mathrm{v}^{\nu}}+\mathrm{v}^{\mu}\right\},$ (B.6)
@@
$m\mathrm{v}^{\nu}\nabla_{\nu}\mathrm{v}^{\mu}=g\tilde{F}^{\mu\nu}\mathrm{v}_{\nu}+\mathcal{F}^{\mu},$ (B.7)
@@
$m\frac{d}{dt}(\gamma\bm{v})=g\bm{B}-\frac{C(qg)^{2}n_{\mathrm{p}}}{4\pi m_{\mathrm{p}}}\frac{\bm{v}}{v}.$ (B.8)
Let us now compare the deceleration and acceleration terms in (B.8) in cosmic environments, by estimating their ratio:
$\epsilon=\frac{1}{gB}\frac{C(qg)^{2}n_{\mathrm{p}}}{4\pi m_{\mathrm{p}}}.$ (B.9)
@@
Let us comment on effects that we have not taken into account. Firstly, the above calculation is based on classical single-particle scatterings. A quantum field theory computation may give rise to corrections to the scattering cross section, especially in the ultrarelativistic regime. We also note that the picture of single-particle scatterings breaks down at distances larger than the Debye length, where plasma effects become important. These have been studied in [47, 48] for slow MMs, for which it was shown that the energy loss from plasma effects is typically not significantly larger than that from single-particle scatterings. Finally, we remark that MMs can further loose energy through ionization of neutral particles, atomic excitations, bremsstrahlung, electron-pair production, and photonuclear interactions; the last two processes become particularly important in the ultrarelativistic regime. (See e.g. [18] for a study of various energy-loss processes for MMs. A rough estimate of the interactions with neutral particles in the ISM is given in [34], according to which they are negligible for a wide range of parameters.) It would be important to study all these effects, in particular in the ultrarelativistic regime. We leave this for future work.
@@
$P\sim g^{2}\gamma^{6}\left\{(\dot{\bm{v}})^{2}-(\bm{v}\times\dot{\bm{v}})^{2}\right\}=\frac{g^{4}B^{2}}{m^{2}},$ (B.10)
where an overdot denotes a time-derivative. Upon moving to the far right-hand side, we have considered a MM accelerating along a homogeneous magnetic field and dropped the cross product, and also used the equation of motion (B.8) without the drag force term. The ratio between the radiative energy loss, and energy gain per time from the acceleration in the magnetic field, $gBv$ (here we take $g>0$), is thus
$\delta\sim\frac{g^{3}B}{m^{2}v}.$ (B.11)
@@
Appendix C Monopole-magnetic field oscillation
If MMs accelerated in magnetic fields do not dissipate their kinetic energy into the ambient plasma, they eventually return the energy to the magnetic fields. In this way the energy can move back and forth between the magnetic field and monopoles (which would be the magnetic analog of the Langmuir oscillation [49]). However one would expect such oscillations to be subject to Landau damping, if the phase velocity of the magnetic field were smaller than the random velocity of the individual MMs [27].
In order to estimate the phase velocity, we start by considering a homogeneous one-dimensional system where the magnetic field has a component $B$ along a certain direction, and the MMs with charge $g$ ($>0$) have a velocity component $v$ along the same direction. (In this appendix, $B$ and $v$ are not the amplitudes and thus can be negative.) Then by combining the MMs’ equation of motion, $m\,d(\gamma v)/dt=gB$, with the conservation of total energy density $\rho_{\mathrm{tot}}=nm\gamma+B^{2}/2$ where $n$ is the MM number density, one obtains an evolution equation for the magnetic field,
$\frac{d^{2}B}{dt^{2}}+\Omega^{2}B=0.$ (C.1)
@@ Appendix C Monopole-magnetic field oscillation
$\Omega=\frac{g}{\gamma^{3/2}}\left(\frac{n}{m}\right)^{1/2}.$ (C.2)
@@
$v_{\mathrm{ph}}\sim 10^{-5}\left(\frac{v_{\mathrm{MW}}}{10^{-3}}\right)^{-1/2}\left(\frac{F_{\mathrm{MW}}}{10^{-15}\,\mathrm{cm}^{-2}\mathrm{sec}^{-1}\mathrm{sr}^{-1}}\right)^{1/2}\left(\frac{g}{g_{\mathrm{D}}}\right)\left(\frac{m}{10^{18}\,\mathrm{GeV}}\right)^{-1/2}\left(\frac{\lambda_{\mathrm{G}}}{1\,\mathrm{kpc}}\right).$ (C.3)
We have rewritten $n$ in terms of the MM flux in the rest frame of the Milky Way Galaxy (cf. below (3.2); however here we consider both clustered and unclustered monopoles). The reference value for the flux is set to the Galactic Parker limit at $m=10^{18}\,\mathrm{GeV}$ (cf. Figure 4). Note that MMs with a Dirac charge and mass $m\gtrsim 10^{18}\,\mathrm{GeV}$ can cluster with the Galaxy [27, 34]. Such clustered MMs obtain a virial velocity of $10^{-3}$, which is larger than the magnetic field’s phase velocity; hence the oscillation is expected to rapidly evaporate. The phase velocity can in principle become larger for smaller masses, however light MMs do not cluster and hence should pass through the Galaxy before (completely) returning the energy to the Galactic fields. A flux much larger than the Parker limit can also increase the phase velocity beyond the virial velocity, however [50] claimed that oscillating Galactic fields do not match with observations, and also that the inhomogeneity in the MM distribution further leads to the damping of the oscillations. For these reasons, one can trust the Parker bound without worrying about the possibility that the Galactic fields survive as oscillating fields. [^12] Primordial magnetic fields can also be used to obtain Parker-type bounds, because MMs effectively dissipate energy into the primordial plasma and thus energy oscillations do not happen even without Landau damping [32, 33, 34]. This is in contrast to the tiny energy loss of MMs in the present-day IGM or ISM, which we showed in Appendix B.
@@
$v_{\mathrm{ph}}\sim 10^{-5}\,v_{\mathrm{CMB}}^{-1/2}\left(\frac{F_{\mathrm{CMB}}}{10^{-27}\,\mathrm{cm}^{-2}\mathrm{sec}^{-1}\mathrm{sr}^{-1}}\right)^{1/2}\left(\frac{g}{g_{\mathrm{D}}}\right)\left(\frac{m}{10^{10}\,\mathrm{GeV}}\right)^{-1/2}\left(\frac{\lambda_{\mathrm{I}}}{1\,\mathrm{Mpc}}\right).$ (C.4)
@@
Appendix D Monopole acceleration in intergalactic and galactic magnetic fields
(a) $B_{\mathrm{I}}=10^{-10}\,\mathrm{G}$.
(b) $B_{\mathrm{I}}=10^{-15}\,\mathrm{G}$.